OpenStudy (anonymous):

For each of the following equations, find the value(s) of the constant α so that the equation has exactly one solution, and determine the solution for each value. x^2+αx+3α+3 To solve i set the discriminant to zero using the following coefficients a=1 b=α c= 3α+3 giving me α^2-4(1)(3α+3) α^2-12α-12 then I plugged that into the quadratic to get the two α's but my online HW is saying that it is incorrect.

ganeshie8 (ganeshie8):

that looks good to me ! must be some syntax issue with the grader

ganeshie8 (ganeshie8):

try again with `sqrt` or ()^(1/2) etc.. http://www.wolframalpha.com/input/?i=solve+%CE%B1%5E2-4%281%29%283%CE%B1%2B3%29+%3D+0

ganeshie8 (ganeshie8):

lol batman, goto sleep :P

ganeshie8 (ganeshie8):

btw, the question is asking for the solutions of given quadratic, NOT the values of \(\alpha \)

OpenStudy (nincompoop):

find the value(s) of the constant α so that the equation has exactly one solution, and determine the solution for each value.

ganeshie8 (ganeshie8):

yeah so try entering the value of quadratic at the \(\alpha \) values

OpenStudy (dan815):

th^

OpenStudy (dan815):

for the solution to have just 1 root, the discriminatnt must be zero

OpenStudy (cp9454):

ans is \[\alpha = 6\pm4\sqrt3\]

ganeshie8 (ganeshie8):

\[\large x = \dfrac{-\alpha\pm 0 }{2}\]

OpenStudy (dan815):

oh u just messed up your coefficient b

OpenStudy (anonymous):

looks like it just wanted me to enter in a very specific syntax, thanks all.

OpenStudy (dan815):

no wait u didnt.... that is wierd

ganeshie8 (ganeshie8):

Ryan, what did u get for \(\alpha \) ?

OpenStudy (cp9454):

lol funny

OpenStudy (dan815):